a) For persons who can only see objects clearly that lie between 17.5 cm and 225 cm from their eyes, what is the change in refractive power of their eyes, in diopters? (Assume a lens-to-retina distance of 2.00 cm.)

    Answer:                P=44.29 for 17.5cm and P = 49.56 for 225cm, and change in P = 5.27 (CORRECT)

b) What strength of contact lenses should be prescribed to correct this myopia (answer in diopters)?

    Answer                 P=-0.444 (CORRECT)

c) If they read a book while wearing contact lenses, what is the minimum distance from their eyes at which they can focus on the book?

    JUST NEED HELP WITH THIS PART
Answer #1

You got for

part a) P1 =5.27D

and for Part b) P2 =-0.444D

Part c)

The minimum distance from their eyes at which they can focus on the book is P =P2-P1 =-0.444-5.27 =-5.714D